# The Impact of Renewable Energy Communities in the Italian Day-Ahead Electricity Market: A Scenario Analysis

### v1

@@ Abstract
This paper evaluates the economic impact of Renewable Energy Communities (RECs) on the Italian wholesale power market. Combining a bottom-up engineering approach with a short-run economic impact assessment, the study begins by mapping existing and emerging RECs in Italy. We identify key characteristics of RECs, such as average installed capacity, institutional profiles of members, types of renewable systems used, and distribution across Italy’s electricity market zones. This mapping yields representative REC configurations, which are employed within a bottom-up engineering model to generate energy injection and self-consumption profiles for different REC prosumer and producer categories (residential, public, small and medium enterprise, non-profit organization, and standalone installation), considering the different levels of solar irradiance in Italy based on latitude. These zonal results, aggregated on an hourly basis, inform the implementation of the synthetic counterfactual approach, which develops alternative scenarios (e.g., 5 GW target for REC-driven capacity set by Italian policy for 2027) to assess the impact of REC-driven injection and self-consumption on the Italian day-ahead power market. The findings suggest that REC deployment can increase equilibrium quantities during daylight in most of the time, while decreasing equilibrium quantities mostly during the cold months, as electrified heating drives greater self-consumption and offsets lower grid injections. Both positive and negative effects on equilibrium quantities suggest that REC deployment also has a potential to reduce wholesale electricity prices. Moreover, by reducing grid exchanges through higher self-consumption, REC proliferation can alleviate pressure on the distribution system.
@@ 1 Introduction
The current integration of Renewable Energy Communities (RECs) in national electricity markets has made it necessary to accurately describe the behaviour of decentralized energy producers and consumers who unite into RECs. Depending on the historical path and national legislation, energy communities could act on the market as utilities do, thereby having the possibility to arbitrate between different market prices. However, RECs typically integrate a small number of users (Lupi et al., 2021). Hence, their direct interaction with wholesale markets is impossible or not allowed by a regulator. In such cases, members of RECs should be supplied by external utilities and be considered as simple users (or clients) of such utilities and/or external aggregators. Nonetheless, these users may exhibit various characteristics, acting as consumers, prosumers, or prosumagers (Sioshansi, 2019).
A comprehensive simulation model that optimizes energy sharing among households and municipal buildings within a REC is provided by Casalicchio et al. (2025), supporting REC planning through sector-coupling and economies of scale. Magni et al. (2025) investigated Italy’s transition from experimental to definitive policies for RECs, by focusing on the economic impacts implied for different RECs configurations across regions. The authors found that the current policy framework reduces regional disparities but may discourage investment in the South of Italy, despite its solar potential. Crucially, understanding the composition of prosumers (users who both generate and consume energy) within RECs (and seasonal heterogeneity) is key to investigate the role of RECs in reshaping market price signals, balancing local demand, and reducing reliance on the grid.
RECs have been deployed in many parts of the world. Europe has the largest number of RECs, with almost 4,000 initiatives and 900,000 members (Koltunov et al., 2023). European RECs deserve particular attention because of the scale of the phenomenon and their formal recognition at the EU level by the RED-II (European Union, 2018) and IEM Directives (European Union, 2019), which triggered their further growth. Outside Europe, RECs exist in the USA, Canada, Costa Rica, Australia and New Zealand, where they are mostly represented by community solar utility initiatives or historic rural cooperatives (Kolesar, 2022). In many states of the Global South such as Brazil and India, RECs are established as off-grid microgrids in areas with limited or no preexisting access to the grid (Hochstetler and Born, 2022; Thapar et al., 2017). Due to the lack of legislation and large heterogeneity of practices, RECs outside Europe remain a niche sector, thus restricting the potential for their deployment (Koltunov and De Vidovich, 2025). Therefore, the impacts of RECs on electricity markets may be negligible, and thus hard to investigate in such jurisdictions.
RECs are undoubtedly considered a unique instrument for a just energy transition. According to the EU legislation, there are two core features that define a REC:
@@
There are several obvious economic and social benefits for citizens willing to join RECs and for local communities hosting them. For example, prosumers and consumers could benefit from a collective generation facility and from shared energy, while local communities usually benefit from the redistribution of profits derived from REC activities (e.g., investment in local infrastructure, nature-based solutions, energy education projects). Economic impacts of RECs can be various, ranging from effects on members and investors, on local and regional economies, and on electricity market and its stakeholders.
Table 1 depicts a taxonomy of economic impacts of RECs. The pre-Directives literature (Bauwens et al., 2016; Brummer, 2018; Brummer et al., 2017; Candelise and Ruggieri, 2017; Heras-Saizarbitoria et al., 2018; Holstenkamp and Kahla, 2016; K.Huntala, 2016; Koltunov and Bisello, 2021; Kooij et al., 2018; Magnani and Osti, 2016; Magnusson and Palm, 2019; Tricarico, 2018; Vernay and Sebi, 2020; Wirth, 2014; Walker et al., 2010; Berka and Creamer, 2018; Moroni et al., 2019) mainly explored micro-scale socio-economic impacts on REC members with a specific focus on domains of management and sociology, while rarely examining the meso [^1] Pre-Directives, there are only few economic studies with those exploring RECs impact on the regional economies using input-output models (Lantz and Tegen, 2011; Okkonen and Lehtonen, 2016; Torgerson, 2006; Phimister and Roberts, 2012; Allan et al., 2011; Bere et al., 2015; Entwistle et al., 2014) and macro effects of RECs on the economy and markets.
@@
The macro impact of REC deployment on electricity systems remains an underexplored domain (Koltunov and De Vidovich, 2025). Robinson and del Guayo (2022) propose two categories: systemic impact and stakeholder impact. The first category explores the impact of RECs on the electricity systems overall, which entails multiple spillovers on the operations of various actors. Examples of such systemic impacts include changes in distribution charges by Distribution System Operators (DSOs), transmission charges by Transmission System Operators (TSOs), distribution and transmission system expansion, wholesale prices, changes in collected taxes and levies embedded in electricity tariffs, and market competitiveness (Berg et al., 2024; Frieß et al., 2025). The second category examines the impact of RECs on specific stakeholder groups, such as non-members, retailers, generators, DSOs, aggregators, and technology providers. The present article belongs to the first category, as we focus on the merit-order effect and the impact of RECs on the day-ahead (DA) power market equilibrium, which entails multiple spillovers on the operations of various stakeholders.
In contrast to the multiple benefits at the organizational and local levels, REC deployment could lead to adverse effects from an economic perspective at macro level: e.g., an increase in distribution grid charges in the presence of volumetric tariffs, which could adversely affect households non-participating in RECs. Post-Directives, several economic studies have appeared on the effect of RECs on the electricity system (Backe et al., 2022; Sarfarazi et al., 2020; Fuentes González et al., 2022; Di Silvestre et al., 2021; Boccard and Goetz, 2025). However, most post-Directive studies focus on the analysis of peer-to-peer trading, an innovative feature of REC that is yet rarely encountered in real-world applications [^2] More details on RECs classifications can be found in Rossetto et al. (2022), Kolesar (2022), Koltunov et al. (2025b) (Castellini et al., 2021; Chen and Gao, 2024; Dong and Li, 2024; Glachant and Rossetto, 2021; Hahnel et al., 2020; Hahnel and Fell, 2022; Nieto-Martin et al., 2019; Sousa et al., 2019).
By complementing existing achievements in the literature, our paper investigates the interactions of Italian RECs with the national wholesale DA market. In terms of business model archetype, Italian RECs can be classified as a ’community collective generation’. [^3] In this type of REC, generation facilities must be connected to the same voltage substation and all members should live in its proximity, while individual members (households/SMEs/public entities/non-profit organisations) maintain their own retailers that take care of their residual demand. This type of REC is the one that is mostly adopted in Italy. This type of REC includes prosumers, consumers, and producers (i.e., the collective generation facility). During periods of excess of generation, RECs supply energy to the grid, thereby increasing renewable dispatch on the supply-side of the wholesale market. At the same time, these RECs aim for higher self-consumption, especially when the selling price is lower than the retail price, thereby decreasing the aggregated load on the demand-side of the market.
For the scope of this paper, we state our research question as follows:
@@
We selected Italy as a case study for several reasons. First, Italian RECs have shown substantial growth in recent years. [^4] In 2023, there were around 50 RECs aligned with new regulation in Italy (Koltunov et al., 2023). In the beginning of 2025, there are nearly 362 RECs either in operational or design phases. Since the transposition of the EU Directives to Italian legislation, 362555Data includes RECs in both operational and design phases as of January 2025 new RECs were constructed. The primary factor behind this notable deployment is the premium tariff, granted at the end of 2023 (MASE, 2023). To date, this incentive is the only available subsidy for new photovoltaic installations in the country. Second, the Italian wholesale day-ahead power market is well-suited for our research because of its transparency regarding hourly supply and demand bids placed by market operators. Third, Italian regulation is rather innovative and can be characterized as a “virtual scheme”, where REC members (prosumers and consumers) do not physically share energy in a microgrid. [^6] In contrast, CEER (2019) reports that “Some Member States, such as France and Austria, have developed a framework for collective self-consumption, where energy can be shared within a group of customers, without requiring direct involvement of a supplier.” Finland adopted a concept of energy sharing that is typically limited to apartment blocks or single housing associations. Similar to Finland, German “collective self-consumers” concept relates largely to occupants of the same building or small groups rather than a rule allowing multi-building or multi-substation virtual sharing across a distribution grid. Instead, they retain their existing retail contracts while sharing energy virtually, for which they are remunerated at the end of the year by the GSE [^7] GSE - in Italian, Gestore del Sistema Energetico - is the energy system manager, a public company that manages incentives in the energy sector on the state’s behalf. (Schiavo et al., 2022). This innovative regulation provides an interesting case for investigating its implications. According to new rules (ARERA, 2022), REC members can benefit in four ways: a premium tariff for shared energy; valorization of avoided grid usage – calculated based on shared energy due to connection to the same primary substation; energy sales; and energy self-consumption (Blasuttigh et al., 2025). Figure 1 summarizes the benefits.
@@
Notably, only the sales of injected energy (by prosumers and producers) and the self-consumed energy (by prosumers) affect the wholesale market, with the former impacting the supply-side and the latter the demand-side. The incentivized ‘shared energy’ is merely a virtual concept that does not directly trigger wholesale market changes. Therefore, we analyze only the operations of REC members that directly influence the market – namely, prosumers and producers – and disregard consumers. [^8] When joining a REC, consumers do not necessarily change their consumption profile in the ‘virtual scheme’ applied in Italy, therefore the wholesale market outcome does not amend. Nonetheless, a REC manager/aggregator might advice or order consumers to increase/decrease consumption in certain hours for the maximization of the common shared energy. Such advanced coordination requires smart infrastructure at the consumers’ premises and could potentially affect the wholesale market situation. However, to the authors’ best knowledge, no RECs in Italy have employed such level of coordination yet. Therefore, this aspect lies out of the scope of the current study and remains for further investigation.
@@
Figure 1: Economic benefit to members of RECs in Italy (excl. explicit capital subsidy).
@@
We apply an innovative mixed-model strategy in our study. In the first stage of analysis, we design a bottom-up engineering model that simulates the hourly behavior of various typical prosumers and producers within a REC. Then, we project the model’s results onto several scenarios based on the actual REC deployment status in 2024 and the policy target for 2027. The outcome of the first stage is then inserted into our synthetic modelling approach, which simulates short-run schedule for the DA market on an hourly basis, from which we derive the effect of RECs on the merit order and aggregated demand.
@@
Hence, the novelty of our work is twofold. First, we provide a theoretical contribution to the understanding of the impact of different types of RECs on the electricity system – specifically, at the wholesale power market level –, where the changes in the equilibrium quantity and price due to REC deployment have multiple spillovers on the operations of other generators, retailers, and final consumers. Notably, heterogeneous REC membership becomes typical in many EU countries, especially after the implementation of the RED-II and IEM Directives. However, the existing literature on modeling RECs and prosumers’ behaviour mostly simulates the impacts of residential prosumers at a single geographical location, whereas our study simulates RECs as a collection of prosumers with very different consumption profiles (public, residential, small and medium enterprises, non-profit organizations) and standalone producers at various climatic locations, which closely reflects the country-scale deployment situation. Our contribution is tested empirically for the entire year 2024 in the DA Italian pool using a synthetic approach to simulate real-world competition among market agents. By undertaking such a detailed modeling exercise, our theoretical contribution may achieve higher external validity, especially for jurisdictions and markets with similar technical definition of a REC (e.g., Spain, Greece, Czech Republic), climate conditions and generation profiles (e.g., Spain, Greece).
@@
Second, we provide a novel mixed-model approach to assess the country-scale impact of prosumers and producers aggregated into a REC on wholesale market scheduling. The first stage of our methodology begins with detailed data collection aimed at deriving available information on REC compositions, generation plant capacities, consumption profiles of different categories of prosumers, and the geographical distribution of RECs. By utilizing real-world data, we design the engineering model to obtain assumed prosumers’ generation and load profiles, which we then project to the policy-targeted deployment level (“Policy Scenario”) and other predicted deployment levels (”Half-way Scenario”, ”Business-as-usual Scenario”). In the second stage, the methodology is expanded by incorporating the previous outputs into an economic synthetic model. In a nutshell, our methodological approach can be used by policymakers and interested stakeholders to determine the actual and predicted systemic effects of REC deployment.
@@
The remainder of the article is organized as follows. In Section 2, we explore the existing literature on systemic impacts of RECs. It then concludes with a review of studies that utilized a methodology similar to ours, i.e. the ”synthetic approach” for modeling electricity markets. Section 3 dives into the methodology employed in this paper, while Section 4 displays the major results from the analysis. Section 5 embeds the results into the academic debate, discusses their implications for policymaking and future research directions. Section 6 concludes with final remarks.

### v2

@@ Abstract
This paper evaluates the economic impact of Renewable Energy Communities (RECs) on the Italian wholesale power market. Combining a bottom-up engineering approach with a short-run economic impact assessment, the study begins by mapping existing and emerging RECs in Italy. We identify key characteristics of RECs, such as average installed capacity, institutional profiles of members, types of renewable systems used, and distribuiton across Italy’s electricity market zones. This mapping yields representative REC configurations, which are employed within a bottom-up engineering model to generate energy injection and self-consumption profiles for different REC prosumer and producer categories (residential, public, small and medium enterprise, non-profit organization, and standalone installation), considering the different levels of solar irradiance in Italy based on latitude. These zonal results, aggregated on an hourly basis, inform the implementation of the synthetic counterfactual approach, which develops alternative scenarios (e.g., 5 GW target for REC-driven capacity set by Italian policy for 2027) to assess the impact of REC-driven injection and self-consumption on the Italian day-ahead power market. The findings suggest that the REC deployment can increase equilibrium quantities during daylight over most of the time, while decreasing equilibrium quantities mostly during the cold months. The differing load profiles of institutional REC members, compared to residential members, drive most of the decreasing effect. Both positive and negative effects on equilibrium quantities suggest that REC deployment also has a potential to reduce wholesale electricity prices. Moreover, when reducing grid exchanges, REC proliferation can alleviate pressure on the transmission system.
@@ 1 Introduction
The current integration of Renewable Energy Communities (RECs) in national electricity systems has made it necessary to accurately describe its market consequences. Depending on the historical path and national legislation, energy communities could act on the market as utilities do, thereby having the opportunity to arbitrate between different market prices. However, RECs typically integrate a small number of users (Lupi et al., 2021). Hence, their direct interaction with wholesale markets is impossible or not allowed by a regulator. In such cases, members of RECs should be supplied by external utilities and be considered as simple users (or clients) of such utilities and/or external aggregators. Nevertheless, these users may exhibit various characteristics, acting as consumers, prosumers, or prosumagers (Sioshansi, 2019), thus implicitly affecting electricity markets. Understanding the composition and behaviour of prosumers within RECs—including their seasonal and geographical heterogeneity—is essential for assessing how these configurations reshape wholesale market price signals and reduce user dependence on the grid. Such insights are critical for designing policies that generate benefits not only for REC participants but also for the broader energy system.
RECs are undoubtedly considered a unique instrument for a just energy transition. There are two core features that define a REC:
@@
Economic impacts of RECs can be various, ranging from effects on members and investors, on local and regional economies, and on electricity market and its stakeholders. Table 1 depicts a taxonomy of economic impacts of RECs. The pre-Directives literature (Bauwens et al., 2016; Brummer, 2018; Brummer et al., 2017; Candelise and Ruggieri, 2017; Heras-Saizarbitoria et al., 2018; Holstenkamp and Kahla, 2016; K.Huntala, 2016; Koltunov and Bisello, 2021; Kooij et al., 2018; Magnani and Osti, 2016; Magnusson and Palm, 2019; Tricarico, 2018; Vernay and Sebi, 2020; Wirth, 2014; Walker et al., 2010; Berka and Creamer, 2018; Moroni et al., 2019) mainly explored micro-scale socio-economic impacts on REC members with a specific focus on domains of management and sociology, while rarely examining the meso [^1] Pre-Directives, there are only few economic studies with those exploring RECs impact on the regional economies using input-output models (Lantz and Tegen, 2011; Okkonen and Lehtonen, 2016; Torgerson, 2006; Phimister and Roberts, 2012; Allan et al., 2011; Bere et al., 2015; Entwistle et al., 2014) and macro effects of RECs on the economy and markets.
In contrast to the multiple benefits at the organizational and local levels, REC deployment could lead to adverse effects from an economic perspective at macro level. For example, the extra injection of intermittent renewable energy could cause higher ramp/start events of expensive peaking plants, an increase in distribution grid charges due to RECs in the presence of volumetric tariffs could adversely affect households non-participating in RECs. Post-Directives, several economic empirical studies have appeared on the effect of RECs on the electricity system (Backe et al., 2022; Sarfarazi et al., 2020; Fuentes González et al., 2022; Di Silvestre et al., 2021; Boccard and Goetz, 2025). However, most post-Directive studies focus on the analysis of peer-to-peer trading, an innovative feature of REC that is yet rarely encountered in real-world applications [^2] More details on RECs classifications can be found in Rossetto et al. (2022), Kolesar (2022), Koltunov et al. (2025b) (Castellini et al., 2021; Chen and Gao, 2024; Dong and Li, 2024; Glachant and Rossetto, 2021; Hahnel et al., 2020; Hahnel and Fell, 2022; Nieto-Martin et al., 2019; Sousa et al., 2019).
@@
The macro impact of REC deployment on electricity systems remains an underexplored domain (Koltunov and De Vidovich, 2025). Robinson and del Guayo (2022) propose two categories: systemic impact and stakeholder impact. The first category explores the impact of RECs on the electricity systems overall, which entails multiple spillovers on the operations of various actors. Examples of such systemic impacts include changes in distribution charges by Distribution System Operators (DSOs) (Berg et al., 2024), transmission charges by Transmission System Operators (TSOs), distribution and transmission system expansion, wholesale prices, changes in collected taxes and levies embedded in electricity tariffs, and market competitiveness. The second category examines the impact of RECs on specific stakeholder groups, such as non-members, retailers, generators, DSOs, aggregators, and technology providers. The present article belongs to the first category, as we focus on the merit-order effect and the impact of RECs on the Italian day-ahead (DA) power market equilibrium, which entails multiple spillovers on the operations of various stakeholders.
In terms of business model archetype, Italian RECs can be classified as a ’community collective generation’. [^3] In this type of REC, generation facilities must be connected to the same voltage substation and all members should live in its proximity, while individual members (households/SMEs/public entities/non-profit organisations) maintain their own retailers that take care of their residual demand. This type of REC is the one that is mostly adopted in Italy. This type of REC includes prosumers, consumers, and producers (i.e., the collective generation facility). During periods of excess of generation, RECs supply energy to the grid, thereby increasing renewable dispatch on the supply-side of the wholesale market. At the same time, these RECs aim for higher self-consumption, especially when the selling price is lower than the retail price, thereby decreasing the aggregated load on the demand-side of the market. For the scope of this paper, we state our research question as follows:
@@
Europe has the largest number of RECs, with almost 4,000 initiatives and 900,000 members (Koltunov et al., 2023). European RECs deserve particular attention because of the scale of the phenomenon and their formal recognition at the EU level by the RED-II (European Union, 2018) and IEM Directives (European Union, 2019), which triggered their further growth. Due to the lack of legislation and large heterogeneity of practices, RECs outside Europe remain a niche sector, thus restricting the potential for their deployment (Koltunov and De Vidovich, 2025). Within Europe, we selected Italian RECs as a case study for several reasons. First, they have shown substantial growth during last few years. [^4] In 2023, there were only 50 RECs aligned with new regulation in Italy (Koltunov et al., 2023). Currently, there are around 362555Data includes RECs in both operational and design phases as of January 2025 new RECs in Italy and the number continues to grow. The primary factors behind this notable deployment include the nuanced policy that allowed institutional actors to participate in RECs [^6] Historically, RECs were considered citizen-driven initiatives, which limited their deployment. and the premium tariff for plants of up to 1 MW capacity within a REC configuration (MASE, 2023). To date, the incentive for RECs is the only available subsidy for new photovoltaic installations in the country. Italy allocated €5.7 billion to support 5GW REC deployment: €3.5 billion for premium tariffs (financed via a levy on electricity consumption) and €2.2 billion from the National Recovery and Resilience Fund to cover up to 40% of CAPEX in small municipalities. These figures highlight the strategic importance of assessing the effectiveness and system-wide value of RECs. Second, the Italian wholesale day-ahead power market is well-suited for our research because of its transparency regarding hourly supply and demand bids placed by market operators. Third, Italian regulation is rather innovative and can be characterized as a “virtual scheme”, where REC members (prosumers and consumers) do not physically share energy in a microgrid. [^7] In contrast, CEER (2019) reports that “Some Member States, such as France and Austria, have developed a framework for collective self-consumption, where energy can be shared within a group of customers, without requiring direct involvement of a supplier.” Finland adopted a concept of energy sharing that is typically limited to apartment blocks or single housing associations. Similar to Finland, German “collective self-consumers” concept relates largely to occupants of the same building or small groups rather than a rule allowing multi-building or multi-substation virtual sharing across a distribution grid. Instead, they retain their existing retail contracts while sharing energy virtually, for which they are remunerated at the end of the year by the GSE [^8] GSE - in Italian, Gestore del Sistema Energetico - is the energy system manager, a public company that manages incentives in the energy sector on the state’s behalf. (Schiavo et al., 2022). This innovative regulation provides an interesting case for investigating its implications for the DA wholesale market. According to new rules (ARERA, 2022), REC members can benefit in four ways: energy self-consumption; a premium tariff for shared energy; valorization of avoided grid usage – calculated based on shared energy due to connection to the same primary substation; and energy sales (Blasuttigh et al., 2025). Figure 1 illustrates the benefits while Appendix 7.2 explains the formulas behind cashback components.
@@
Notably, only the sales of injected energy (by prosumers and producers) and the self-consumed energy (by prosumers) affect the wholesale market, with the former impacting the supply-side and the latter the demand-side. The incentivized ‘shared energy’ is merely a virtual concept that does not directly trigger wholesale market changes. Therefore, we analyze only the operations of REC members that directly influence the market – namely, prosumers and producers – and disregard consumers. [^9] When joining a REC, consumers do not necessarily change their consumption profile in the ‘virtual scheme’ applied in Italy, therefore the wholesale market outcome does not amend because of them. Nonetheless, a REC manager/aggregator might advice consumers to change consumption in certain hours for the maximization of the common shared energy. Such advanced coordination requires smart infrastructure at the consumers’ premises and could potentially affect the wholesale market situation. However, to the authors’ best knowledge, no RECs in Italy have yet employed such level of coordination. Therefore, this aspect lies beyond the scope of the current study and remains for further investigation. From this perspective, our study pertains to the literature on the impact of prosumers on wholesale markets, although using Italian RECs as a specific case.
@@
Figure 1: A net cash flow for REC members in Italy (excl. explicit capital subsidy).
@@
We apply an innovative mixed-model methodology. In the first stage of analysis, we design a bottom-up engineering model that simulates the hourly behavior of various typical prosumers and producers within a REC. Then, we project the model’s results onto several scenarios, which are based on the policy target for 2027 and on an actual REC deployment status in 2024. The outcome of the first stage is then inserted into our synthetic model that simulates short-run schedule for the DA market on an hourly basis, from which we derive the effect of RECs on the merit order and aggregated demand.
@@
Hence, the novelty of our work is twofold. First, we provide a theoretical contribution to the understanding of the seasonal and hourly impact of prosumers with different load profiles on the wholesale power market. We take Italian RECs as a specific case study. The heterogeneous membership within RECs becomes typical in many EU countries, especially after the implementation of the RED-II and IEM Directives. However, the existing literature on modeling prosumers’ behaviour mostly simulates the impacts of residential prosumers at a single geographical location, whereas our study simulates RECs as a collection of prosumers with very different consumption profiles (public, residential, small and medium enterprises, non-profit organizations) and standalone producers at various climatic locations, which closely reflects the country-scale REC deployment situation. Our contribution is tested empirically for the entire year 2024 in the DA Italian pool using a synthetic approach to simulate real-world competition among market agents. By undertaking such a detailed modeling exercise, our theoretical contribution may achieve higher external validity, especially for jurisdictions and markets with similar technical definition of a REC (e.g., Spain, Greece, Czech Republic), climate conditions and generation profiles (e.g., Spain, Greece).
@@
Second, we provide a novel mixed-model approach to assess the country-scale impact of prosumers and producers aggregated into a REC on wholesale market scheduling. The first stage of our methodology begins with detailed data collection, which incorporates information on the composition of prosumers within RECs, the capacities of generating plants, the load profiles of different categories of prosumers, and the geographical distribution of RECs. Using these real-world parameters, we design the engineering model to obtain the assumed prosumers’ generation and load profiles, which we then project to the policy-targeted deployment level and other predicted deployment levels. In the second stage, the methodology is expanded by incorporating the previous outputs into an economic synthetic model. In essence, our methodological approach can be used by policymakers and stakeholders to determine the actual and predicted effects of the growing presence of heterogeneous prosumers, including, but not limited to, those arranged within RECs, on wholesale markets.
@@
The remainder of the article is organized as follows. In Section 2, we explore the existing literature on systemic impacts of RECs and studies that utilized the ”synthetic approach” to modeling electricity markets. Section 3 explains the novel methodology employed in this paper, while Section 4 displays the major results. Section 5 incorporates the results into the academic debate. Section 6 concludes with final remarks and policy recommendations.